# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ VI-D Feature attributions for ETT dataset
VI-E Performance analysis of EmSHAP under different model structure
In this subsection, the performance of EmSHAP under different model structure is compared and analyzed.
VI-E1 Ablation study
An ablation study is performed to reflect the performance of EmSHAP on different standardized models, which involves the combinations of vanilla EBM, the GRU network, and the dynamic masking mechanism.
Table V presents the feature attribution performance and runtime under different cases. As is shown in Table V, although EmSHAP requires more time for training, it demonstrates superior performance in terms of SIC AUC metric. More specifically, the vanilla EBM yields the worst performance, by gradually adding the dynamic masking mechanism and GRU network, the performance increases. This clearly shows that the combined effect of these mechanisms contributes to the better performance of EmSHAP.
TABLE V: Results of ablation experiments
Dataset Explainer SIC AUC Run time Memory usage
@@ VI-E1 Ablation study
MNIST Vanilla EBM 0.883 0.217 $\mathbf{1568.18}$s $\mathbf{1.02}$s $\mathbf{1354.19}$MB $\mathbf{1200.41}$MB
TREC Vanilla EBM 0.808 0.213 $\mathbf{94.88}$s $\mathbf{0.14}$s $\mathbf{995.57MB}$ $\mathbf{923.68}$MB
Vanilla EBM+Dynamic masking 0.814 0.196 102.00s 0.15s 1055.01MB 989.62MB
Vanilla EBM+GRU coupling 0.922 0.142 156.33s 0.17s 1070.90MB 990.03MB
EmSHAP $\mathbf{0.948}$ $\mathbf{0.103}$ 162.75s 0.19s 1085.62MB 994.82MB
ADTI Vanilla EBM 0.863 0.544 $\mathbf{173.22}$s $\mathbf{0.81}$s $\mathbf{889.21}$MB $\mathbf{850.41}$MB
EmSHAP $\mathbf{0.933}$ $\mathbf{0.411}$ 270.90s 0.87s 966.58MB 946.81MB
ETT Vanilla EBM 0.864 0.440 $\mathbf{14.19}$s $\mathbf{1E-3}$s $\mathbf{515.72}$MB $\mathbf{510.33}$MB
Vanilla EBM+GRU couping 0.903 0.292 15.99s 4E-3s 553.37MB 540.10MB
EmSHAP $\mathbf{0.915}$ $\mathbf{0.243}$ 16.55s 4E-3s 554.11MB 540.46MB
VI-E2 GRU network vs LSTM network
Table V shows that inclusion of GRU as the proposal network enhances the performance of EmSHAP. Since LSTM has a similar structure as GRU and might better capture long-term dependencies between variables, it makes sense to see how EmSHAP performs by replacing the GRU network with the LSTM network. For a fair comparison, the sizes of hidden layer for both LSTM and GRU are set as 32. Table VI summarizes the SIC AUC metric, runtime and memory using under the two cases. Table VI shows that using GRU as the proposal network yields more satisfactory results. Moreover, due to its simpler structure, GRU requires less runtime and memory usage. The fewer parameters and simpler structure of GRU reduce the risk of overfitting, which may explain why the SIC AUC results of GRU surpass those of LSTM.
TABLE VI: Proposal network in EmSHAP: LSTM vs GRU
Dataset Proposal network SIC AUC Run time Memory usage
@@ VI-E2 GRU network vs LSTM network
MNIST LSTM 0.952 0.108 3406.11s 2.01s 2067.50MB 1999.63MB
GRU 0.965 0.037 2977.10s 1.63s 1879.91MB 1823.68MB
TREC LSTM 0.943 0.072 265.34s 0.22s 1201.12MB 1190.92MB
GRU 0.948 0.103 162.75s 0.19s 1085.62MB 994.82MB
ADTI LSTM 0.930 0.416 303.00s 0.91s 995.50MB 932.71MB
GRU 0.933 0.411 270.90s 0.87s 966.58MB 946.81MB
ETT LSTM 0.906 0.292 17.39s 6E-3s 565.50MB 552.71MB
GRU 0.915 0.243 16.55s 4E-3s 554.11MB 540.46MB
VI-E3 Dynamic masking vs fixed-rate masking
As an important component of EmSHAP, the dynamic masking mechanism enabling the model performance. To better show the functionality of dynamic masking, the performance of EmSHAP under dynamic masking and fix-rate masking is compared and the results are shown in Table VII.
TABLE VII: SIC AUC results for different masking rate on four datasets
Dataset SIC AUC Masking rate
0.5 0.6 0.7 0.8 0.9 Dynamic
masking
@@ VI-E3 Dynamic masking vs fixed-rate masking
As shown in Fig. VII, the dynamic masking achieves better performance than fixed-rate masking for all cases. In addition, using the fixed-rate masking requires manually tuning the optimal masking rate, which is undesired in practice. In contrast, the dynamic masking mechanism allows the model to learn masking rate automatically, which leads to the better performance [55].
VII Conclusion
This paper proposes an energy-based model for Shapley value estimation (EmSHAP) for feature attribution in deep learning-based model. In EmSHAP, an energy-based model (EBM) is used to estimate the Shapley contribution function under different feature combinations. The EBM consists of a Resnet structure for approximating the energy function and a GRU network for estimating the partition distribution. By mapping the input features into a latent space using the GRU network, the influence of feature ordering to calculation of contribution function can be eliminated. In order to explore a wider range of feature combinations, a dynamic masking mechanism is introduced during training, which empowers EmSHAP with better estimation accuracy and generalization capability. In order to evaluate the performance of the proposed method, detailed theoretical analysis on error bound and extensive application studies are conducted, which shows that EmSHAP has tighter error bound and exhibits higher accuracy compared with state-of-the-art approaches. It should be noted that despite the theoretically universal approximation capability, EmSHAP may not be suitable to handle the small-sample problem. In addition, although EmSHAP exhibits better performance under conditions like noises, outliers, and adversarial perturbations, it still lacks an inherent mechanism to deal with these cases. Future work will focus on addressing these limitations to enhance the robustness and scalability.
Appendix
In this appendix, the error bounds of KernerlSHAP and VAEAC are analyzed.
VII-A Error Bound of KernelSHAP
In KernelSHAP, the calculation of Shapley value is achieved using an additive model, with the contribution function $v(S)$ of the variable set $S$ approximated as follows.
$\small\hat{v}(S)=\beta_{0}+\sum_{i\in S}\beta_{i},$ (A1)
where $\beta_{0}$ is the expectation of the deep learning model outputs, $\beta_{i}$ is the Shapley value of the $i$-th variable. KernelSHAP uses a linear regression model to obtain the Shapley value of each variable, which is estimated by solving the weighted least squares problem as follows.
$\small\mathop{\min}_{\beta_{0},...,\beta_{|D|}}\sum_{S\in D}\psi(S)\left(\hat{v}(S)-v(S)\right)^{2},$ (A2)
Here $\psi(S)\in[0,1]$ is called the Shapley kernel. To solve the above optimization problem, Covert et al. [12] introduced a binary vector $\mathbf{b}\in\mathbb{R}^{|D|+1}$ to determine whether the $i$-th feature is included in the feature subset $S$. Specifically, $b_{i}=1$ indicates the inclusion of $i$-th feature in $S$, while $b_{i}=0$ represents its exclusion. Thus, Eq.(A2) can be rewritten as follows.
Here, $\bm{\beta}=[\beta_{0},...,\beta_{|D|}]^{T}\in\mathbb{R}^{|D|+1}$. The matrix form of Eq.(A3) is as follows.
where $K$ is the number of sampling times, $\mathbf{B}\in\mathbb{R}^{(|D|+1)\times K}$ is a binary matrix indicating which variables are included in $S$, $\bm{\psi}\in\mathbb{R}^{K\times K}$ is a diagonal matrix with Shapley kernel in its diagonal, $\mathbf{v}\in\mathbb{R}^{K\times 1}$ is the contribution function vector.
The optimal estimation $\hat{\bm{\beta}}$ of Eq.(A4) can be obtained by taking the first-order derivative of $\bm{\beta}$ in Eq.(A4) and setting it to zero.
For more details, interested readers can refer to Refs. [11, 12].
Next we will analyze the error bound of KernelSHAP. Let $\hat{\bm{\Sigma}}=\frac{1}{K}\mathbf{B}\bm{\psi}\mathbf{B}^{T}$ and $\hat{\bm{\Gamma}}=\frac{1}{K}\mathbf{B}^{T}\bm{\psi}\mathbf{v}$, $\mathbb{E}[\hat{\bm{\Sigma}}]=\bm{\Sigma}^{*}$ and $\mathbb{E}[\hat{\bm{\Gamma}}]=\bm{\Gamma}^{*}$. Assuming that $\hat{\bm{\Sigma}}$ and $\bm{\Sigma}^{*}$ are invertible, we have $\hat{\bm{\beta}}=\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}$, $\bm{\beta}^{*}=\mathbb{E}[\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}]$. The error bound of KernelSHAP can be described as the mean absolute deviation between the estimated contribution function $\hat{v}(S)$ and the true contribution function $v(S)$, which is related to $\hat{\bm{\beta}}$ and $\bm{\beta}^{*}$ as follows.
$\small\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]\leq\mathbb{E}\left[||\mathbf{b}^{T}||\cdot||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]=(|S|+1)\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right].$ (A6)
The decomposition of $\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]$ is given by,
$\quad\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\mathbb{E}[\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}]||]$ (A7)
$=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\mathbb{E}[\hat{\bm{\Sigma}}^{-1}]\mathbb{E}[\hat{\bm{\Gamma}}]-\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||]$
$=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\bm{\Sigma}^{*^{-1}}\bm{\Gamma}^{*}-\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||],$
where |||||$\cdot$||| | ⋅ | | is the $l_{2}$ norm and $\text{Cov}(\cdot,\cdot)$ is the covariance. Next we introduce three corollaries and show that $||\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||\leq\epsilon_{3}^{2}$ holds in some specific cases, with $\epsilon_{3}$ being a small positive number.
Corollary A1.
(Concentration of $\hat{\bm{\Sigma}}$) For any $\epsilon_{3}>0$, the following holds,
$\small\mathbb{P}(||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\geq\epsilon_{3})\leq 2(|D|+1)^{2}\exp(-2K\epsilon_{3}^{2}).$ (A8)
Proof. Noting that the elements in $\mathbf{B}$ are 0 or 1 and the elements in $\bm{\psi}$ is bounded within $[0,1]$. It can be seen that the elements in $\hat{\bm{\Sigma}}$ and $\bm{\Sigma}^{*}$ are also bounded within $[0,1]$. According to Hoeffdings inequality, each element in $\hat{\bm{\Sigma}}$ is bounded with their expectation values, that is, for any $\epsilon_{3}>0$
By using Booles inequality [56], Eq.(A8) can be obtained.
Corollary A2.
(Concentration of $\hat{\bm{\Gamma}}$)For any $\epsilon_{3}>0$, the following inequality holds.
$\small\mathbb{P}(||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||\geq\epsilon_{3})\leq 2(|D|+1)\exp(-2K\epsilon_{3}^{2}).$ (A10)
Proof. According to Assumption 4, the contribution value of any subset $S$ is bounded within $[0,1]$, that is $v(S)\in[0,1]$. Then $\mathbf{v}$ is also bounded within $[0,1]$ and $\bm{\Gamma}^{*}\in[0,1]$. By using Hoeffdings inequality, each element in $\hat{\bm{\Gamma}}$ is bounded with their expectation values.
By using Booles inequality [56], the concentration of $\hat{\bm{\Gamma}}$ in Eq.(A10) can be obtained.
Corollary A3.
(Concentration of $\hat{\bm{\Sigma}}^{-1}$) For any $\epsilon_{3}>0$, we have,
$\small\mathbb{P}(||\hat{\bm{\Sigma}}^{-1}-{\bm{\Sigma}^{*}}^{-1}||\geq\epsilon_{3})\leq 2(|D|+1)^{2}\exp(-2K\epsilon_{3}^{2}).$ (A12)
Corollary A4.
For a fixed input dimension $|D|$ and any small positive number $\epsilon_{3}$, there exists a large $K$ that
$\small\mathbb{P}\left(||\mathrm{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||\geq\epsilon^{2}_{3}\right)\leq 2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2}),$ (A13)
where $\epsilon_{3}\in\left[\right.\sqrt{\frac{1}{2K}\log(2((|D|+1)^{2}+(|D|+1)))},+\infty\left.\right)$.
Proof. $\left|\left|\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})\right|\right|$ can be decomposed as follows.
$\small\quad\left|\left|\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})\right|\right|=\left|\left|\mathbb{E}\left[\left(\hat{\bm{\Sigma}}^{-1}-\mathbb{E}[\hat{\bm{\Sigma}}^{-1}]\right)\left(\hat{\bm{\Gamma}}-\mathbb{E}[\hat{\bm{\Gamma}}]\right)\right]\right|\right|.$ (A14)
According to Corollaries A2, A3, and Booles inequality, Eq.(A13) can be obtained. Since the probability $\mathbb{P}\left(||\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||\geq\epsilon_{3}\right)\in[0,1]$, the range of $\epsilon_{3}$ is $\left[\right.\sqrt{\frac{1}{2K}\log(2((|D|+1)^{2}+(|D|+1)))},+\infty\left.\right)$.
It should be noted that for a small positive $\epsilon_{3}$, when $K$ is large, the probability $1-2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2})$ is close to 1, indicating that $\left|\left|\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})\right|\right|\leq\epsilon_{3}^{2}$ holds true under almost any conditions.
Thus, Eq.(A7) can be rewritten as follows,
$\quad\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]$ (A15)
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\bm{\Sigma}^{*^{-1}}\bm{\Gamma}^{*}||]+\mathbb{E}[||-\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||]$
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}(\hat{\bm{\Gamma}}-\bm{\Gamma}^{*})||]+\mathbb{E}[||(\hat{\bm{\Sigma}}^{-1}-\bm{\Sigma}^{*^{-1}})\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}$
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}||\cdot||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||]+\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}-\bm{\Sigma}^{*^{-1}}||\cdot||\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}$
$=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}(\hat{\bm{\Gamma}}-\bm{\Gamma}^{*})||]+\mathbb{E}[||\bm{\Sigma}^{*^{-1}}(\hat{\bm{\Sigma}}-\bm{\Sigma}^{*})\hat{\bm{\Sigma}}^{-1}||\cdot||\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}$
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}||\cdot||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||]+\mathbb{E}[||\bm{\Sigma}^{*^{-1}}||\cdot||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\cdot||\hat{\bm{\Sigma}}^{-1}||\cdot||\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}.$
By calculating the upper bounds of the decomposed elements in Eq.(A15) separately and merging them, the upper bound of $\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ can be obtained. Notice that Ref.[57] gives the theoretical analysis of LIME, we follow their contribution and derive the error bound of KernelSHAP. Here, the explicit expressions of $\bm{\Sigma}^{*}$ and its inverse form $\bm{\Sigma}^{*}$ are first taken into consideration. According to Ref. [12], $\bm{\Sigma}^{*}$ has the following form,
that is, the diagonal of $\bm{\Sigma}^{*}$ is $\frac{1}{2}$, and the off-diagonal entries $\alpha$ are computed as follows.
Lemma A1.
where $\mathbf{I}\in\mathbb{R}^{(|D|+1)\times(|D|+1)}$ is an identity matrix and $\mathbf{J}\in\mathbb{R}^{(|D|+1)\times(|D|+1)}$ is a matrix with all ones. The inverse of $\bm{\Sigma}^{*}$ can be denoted as $\bm{\Sigma}^{*{-1}}=\kappa\mathbf{I}+\omega\mathbf{J}$.
By sorting out the above formula,
Thus, the following holds

### v4

@@ VII-A Error Bound of KernelSHAP
Thus, Eq.(A7) can be rewritten as follows,
$\qquad\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ (A16)
@@
[…]
By calculating the upper bounds of the decomposed elements in Eq.(A16) separately and merging them, the upper bound of $\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ can be obtained. Notice that Ref.[73] gives the theoretical analysis of LIME, we follow their contribution and derive the error bound of KernelSHAP. Here, the explicit expressions of $\bm{\Sigma}^{*}$ and its inverse form $\bm{\Sigma}^{*}$ are first into consideration. According to Ref. [14], $\bm{\Sigma}^{*}$ has the following form,
that is, the diagonal of $\bm{\Sigma}^{*}$ is $\frac{1}{2}$, and the off-diagonal entries are $\alpha$, where
Lemma A1.
where $\mathbf{I}\in\mathbb{R}^{(|D|+1)\times(|D|+1)}$ is a identity matrix and $\mathbf{J}\in\mathbb{R}^{(|D|+1)\times(|D|+1)}$ is a matrix with all ones. Let $\bm{\Sigma}^{*{-1}}=\kappa\mathbf{I}+\omega\mathbf{J}$, $\kappa$ and $\omega$ are need to be determined, we have,
By sorting out the above formula,
Thus, we can get,
$\small\begin{cases}\begin{aligned} &(\frac{1}{2}-\alpha)\kappa=1\\ &(\frac{1}{2}-\alpha+K\alpha)\omega+\kappa\alpha=0.\end{aligned}\end{cases}$ (A23)
By solving Eq.(A23), we have,
and
Proposition A1.
(The upper bound of $\bm{\Sigma}^{*{-1}}||$) The $L_{2}$ norm of $\bm{\Sigma}^{*^{-1}}$ is upper bounded with,
[…]
By taking the root square on both sides of Eq.(A26), we can get Eq.(A27).
Similarly, the upper bound of $\bm{\Gamma}^{*}||$ can be obtained in the same way.
Proposition A2.
(Upper bound of $\bm{\Gamma}^{*}||$) The $L_{2}$ norm of $\bm{\Gamma}^{*}$ is upper bounded with,
[…]
Proof. The square $L_{2}$ norm of $\bm{\Gamma}^{*}$ is as follows.
[…]
Since each element in $\bm{\Gamma}^{*}$ is bounded within $[0,1]$, the square $L_{2}$ norm of $\bm{\Gamma}^{*}$ is smaller then its dimension value, that is, $\bm{\Gamma}^{*}||^{2}\leq|D|+1$.
Proposition A3.
[…]
Particularly, $\hat{\bm{\Sigma}}$ is concentrated on $\bm{\Sigma}^{*}$, by using Weyl inequality [74],
[…]
and,
[…]
which can be deduced to
[…]
Furthermore,
[…]
Proof. (Proof of Theorem 2) For fixed feature dimension $|D|$, a small positive $\epsilon_{3}$, and a large sampling number $K$, the final result of $\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ with the probability greater than $1-2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2})$ is as follows.
$\qquad\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$
@@
It can be seen from Eq.(VII-A) that as the data dimension $|D|$ increases, the error estimated by KernelSHAP will increase. On the other hand, the estimation error will decrease as the number of samples $K$ increases.
VII-B Error Bound of Variational autoencoder
Variational autoencoder (VAE) is a typical generative model, which uses a latent variable model to give a probabilistic representation of the unknown data distribution $p(\mathbf{x})$. The basic idea of the variational autoencoder is to generate the observed variable $\mathbf{x}$ through a simple distributed latent variable $\mathbf{z}$. Specifically, VAE first models the joint distribution,
$\small p(\mathbf{x},\mathbf{z})=p(\mathbf{z})p_{\theta}(\mathbf{x}|\mathbf{z}),$ (B1)
where the conditional distribution $p_{\theta}(\mathbf{x}|\mathbf{z})$ is called an encoder, which is modeled by a deep neural network with parameter $\theta$$\mathbf{z}$ is the latent variable with a simple distribution, such as the standard normal distribution. By marginalizing the latent variable, the data distribution can be approximated.
$\small p(\mathbf{x})=\int p(\mathbf{x},\mathbf{z})d\mathbf{z}.$ (B2)
This latent variable model helps to get $p(\mathbf{x})$ easily. However, the calculation of the posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{x})$ during training becomes intractable. To address this challenge, VAE constructs a distribution $q_{\tau}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\theta}(\mathbf{z}|\mathbf{x})$, which is called the decoder parametered with $\tau$. The data log-likelihood of VAE can be expressed as follows.
$\log p_{\theta}(\mathbf{x})$ $=\log p_{\theta}(\mathbf{x})\int q_{\tau}(\mathbf{z}|\mathbf{x})d\mathbf{z}$ (B3)
$=\int q_{\tau}(\mathbf{z}|\mathbf{x})\log p_{\theta}(\mathbf{x})d\mathbf{z}$
$=\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x})}\log p_{\theta}(\mathbf{x})$
$=\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x})}\log\frac{p_{\theta}(\mathbf{x},\mathbf{z})}{p_{\theta}(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x})}\log\frac{p_{\theta}(\mathbf{x},\mathbf{z})q_{\tau}(\mathbf{z}|\mathbf{x})}{p_{\theta}(\mathbf{z}|\mathbf{x})q_{\tau}(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x})}\log\frac{p_{\theta}(\mathbf{x},\mathbf{z})}{q_{\tau}(\mathbf{z}|\mathbf{x})}+\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x})}\log\frac{q_{\tau}(\mathbf{z}|\mathbf{x})}{p_{\theta}(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x})}\log\frac{p_{\theta}(\mathbf{x},\mathbf{z})}{q_{\tau}(\mathbf{z}|\mathbf{x})}+D_{KL}\left(q_{\tau}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}|\mathbf{x})\right),$
where $\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x})}\log\frac{p_{\theta}(\mathbf{x},\mathbf{z})}{q_{\tau}(\mathbf{z}|\mathbf{x})}$ is called evidence lower bound (ELBO) and $D_{KL}(\cdot||\cdot)$ is Kullback-Leibler (KL) divergence. ELBO is the objective function that needs to be optimized by maximizing ELBO via gradient descent, and a fine-trained VAE can be obtained. $D_{KL}\left(q_{\tau}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}|\mathbf{x})\right)$ shows the gap between log-likelihood and ELBO, which is exactly the mismatch between the VAE encoder and the posterior[75].
Ref. [18] introduces a Shapley value approximation approach based on VAE. The contribution function of Shapley value is sampled from the conditional probability $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ generated by the VAE, and the sampling method is just like the method proposed in this article. The log-likelihood introduced in Ref. [18] is as follows.
$\qquad\log p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ (B4)
$=\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]+\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}{p_{\theta}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]$
$=\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]+D_{KL}(q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})||p_{\theta}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$
$\geq\mathbb{E}_{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right].$
This VAE-based Shapley value estimation method trains the parameters in the VAE by maximizing the first term in Eq.(B4), and then samples from the generated conditional distribution $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ to calculate the contribution function $v$ in the Shapley value and the Shapley values of each input.
However, a problem arises here, that the generated conditional distribution $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ convergences to the true conditional distribution $p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ if and only if $D_{KL}(q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})||p_{\theta}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$ goes to zeros. Unfortunately, it’s not explicit to ensure this happens, since $q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})$ is modeled as exponential families, i.e. multivariate Gaussian distribution, the posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{x})$ is not always possible to use the exponential families to approximate. Thus, for an unknown data distribution $p(\mathbf{x})$, $D_{KL}(q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})||p_{\theta}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$ is hardly going to zeros, which indicates that there is a inference gap between the true distribution and the generated distribution.
According to Ref. [37], the inference gap of VAE is decomposed into two components, the approximation gap $\log p_{data}(\mathbf{x})-\log p_{\theta^{*}}(\mathbf{x})$ and the amortization gap $\log p_{\theta^{*}}(\mathbf{x})-\log p_{\theta}(\mathbf{x})$, where $\theta^{*}\in\mathop{\text{argmin}}_{\theta}\log p_{\theta}(\mathbf{x})$. Since the analysis of the model error bound in this paper assumes that the model can achieve theoretical optimal performance after a large number of iterative training, the amortized error of VAE here is negligible. We focus on the approximation gap, followed by the above discussion, we have,
$\qquad\log p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})-\log p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ (B5)
$=D_{KL}\left(q_{\tau}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})||p_{\theta}(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})\right).$
Eq.(B5) shows the difference between the log-likelihood of $\log p_{data}(\mathbf{x})$ and $\log p_{\theta^{*}}(\mathbf{x})$, this difference also leads to the probability density difference,
$\small\left|\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*}})}{p_{data}(\mathbf{x}_{\bar{S}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*}})}-1\right|\leq\delta,$ (B6)
where $\delta$ is the difference of the probability density between $p_{data}$ and $p_{\theta^{*}}$. It should be noted that $\delta$ is positive related to $D_{KL}\left(q_{\tau}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}|\mathbf{x})\right)$.
Similar to Assumption 1, assuming that the parameters of the estimated conditional density $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})$ are unbiased estimates of the parameters of the optimal estimated density $p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})$, and for a small positive $\epsilon_{1}$, the probability density ratio is as follows.
$\small\left|\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}-1\right|\leq\epsilon_{1}.$ (B7)
Then we begin the analysis of VAEAC method. Firstly, $\mathbb{E}[|\hat{v}(S)-v(S)|]$ of VAE-based approach can be derived as follows.
$\qquad\mathbb{E}[|\hat{v}(S)-v(S)|]$ (B8)
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{*})-\mathbb{E}_{p_{data}(\mathbf{x}^{*}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}^{*}_{\bar{S}},\mathbf{x}_{S}^{*})\right]\right|\right]$
@@ VII-B Error Bound of Variational autoencoder
$\qquad\left.\left.+\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{*}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}^{*}_{\bar{S}},\mathbf{x}_{S}^{*})\right]\right|\right]$
$\leq\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{*})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]\right|\right]$